Powell Panther Corporation: Income Statements for Year Ending December 31 (Millions of Dollars) 2015 $1,495.0 $1,300.0 1.121.0 1105.0 $374.0 $195.0 33.0 $332.0 $162.0 29.0 $299.0 $133.0 53.2 $79.8 2016 Sales Operating costs excluding depreciation and amortization EBITDA Depreciation and amortization Earnings before interest and taxes (EBIT) 42.0 Interest 33.0 Earnings before taxes (EBT) Taxes (40%) Net income 119.6 $179.4 $161.0 Common dividends $64.0
Powell Panther Corporation: Balance Sheets as of December 31 (Millions of Dollars) 2016 2015 Assets Cash and equivalents Accounts receivable Inventories $24.0 $20.0 215.0 195.0 257.0 234.0 $496.0 $449.0 423.0 325.0 $919.0 $774.0 Total current assets Net plant and equipment Total assets Liabilities and Equity Accounts payable Accruals Notes payable $98.0 $78.0 104.0 30.0 26.0 $253.0 $208.0 299.0 260.0 $552.0 $468.0 324.6 282.0 42.4 24.0 $367.0 $306.0 $919.0 $774.0 125.0 1 Total current liabilities Long-term bonds Total liabilities Common stock Retained earnings Common equity Total liabilities and equity
Write out your answers completely. For example, 25 million should be entered as 25,000,000 a. What was net operating working capital for 2015 and 2016? 2015: 2016: $ b. What was the 2016 free cash flow? c. How would you explain the large increase in 2016 dividends? 1. The large increase in free cash flow from 2015 to 2016 explains the large increase in 2016 dividends. II. The large increase in net income from 2015 to 2016 explains the large increase in 2016 dividends. III. The large increase in EBIT from 2015 to 2016 explains the large increase in 2016 dividends. IV, The large increase in sales from 2015 to 2016 explains the large increase in 2016 dividends. V. The large increase in retained earnings from 2015 to 2016 explains the large increase in 2 016 dividends. Select-
Answer #1

Part A

Net operating working capital = current assets - (current liabilities - notes payable)

2015 = 449000000-(208000000-26000000) = 267000000

2016 = 496000000-(253000000-30000000)=273000000

Part B

Free cash flow = [EBIT(1 – T) + Depreciation] – [Capital expenditure + NOWC]

= ((332000000*(1-40%)+42000000)+((423000000-325000000+42000000)+(273000000-267000000)= 387200000

Part C

Answer is option B

It has to be net income and basis of dividends are distributed.